# Unsocial Intelligence: an Investigation of the Assumptions of AGI Discourse

### v3

@@ Abstract.
1. Introduction
There is no agreed-upon definition of artificial general intelligence (AGI) (Morris et al., 2023; Mitchell, 2024). Yet from influential AI companies (DeepMind, 2022; OpenAI, 2018; Meaker, 2023; Ingram, 2024) and AI researchers (Agüera y Arcas and Norvig, 2023; OpenAI, 2018; Morris et al., 2023; Chollet, 2019; Bubeck et al., 2023) to the increasingly public worry (Future of Life Institute, 2023; Center for AI Safety, [n. d.]) about existential risks (Bostrom, 2014; McLean et al., 2023; Naudé and Barten, 2023; Noy and Uher, 2022), AGI and human-level AI have become one of the dominant ways to imagine the future potential of AI. At the same time, AGI has had its own skeptics (LeCun, 2022; Toews, 2022; Heaven, 2020; Gebru and Torres, 2024) who risk taking the target of their critiques to be more homogenous and congruent than it actually is.
The lack of homogeneity in current conceptions of AGI is not a bug. It is a feature of the underlying topic: what might it mean for machines to have intelligence that rivals human intelligence? Substantive disagreements come with the value-laden character of both intelligence and of AI as a technology. By value-laden, we mean that political, social, and ethical values can and should shape conceptions of intelligence, technology, and their intersection.(Blili-Hamelin and Hancox-Li, 2023) We lack consensus on the definition of intelligence because of conflicting values. The same applies to the current lack of consensus on AGI.
Ruha Benjamin warns that “a narrow definition of what even counts as technology or intelligence” threatens the ability of communities to imagine worlds worth building.(Benjamin, 2024) What forms of intelligence and technology are worth imagining are political and social questions. However, current approaches to AGI risk mistaking these questions for technical questions. We lay the groundwork for examining these potential mistakes by analyzing and classifying contingent assumptions underlying different definitions of AGI. Contending with the fractured worldviews among conceptions of AGI plants the seeds for resisting harmfully narrow conceptions of machine intelligence.
@@ 1. Introduction
Our investigation is not limited to accounts using the exact phrase “artificial general intelligence”. Like Morris et al. 2023, our topic is the long history of treating “[a]chieving human-level “intelligence” as the “north-star goal” of the AI field (Morris et al., 2023), dating at least as far back as the 1955 Dartmouth AI Conference (McCarthy et al., 1955). Discussions of “human-level AI”, “general AI”, or “AI” (such as in the phrase “strong AI” (Searle, 1980)) fall within the scope of our account.
@@
2. Between human intelligence and technology: AGI’s dual value-laden pedigrees
@@
Building on STS scholarship on the values embedded in technology (Winner, 1980), research communities like FAccT, AIES, and CHI have taken a deep interest in the political, social, and ethical values embedded both in AI tools and in the practices that surround AI (Fishman and Hancox-Li, 2022; Dotan and Milli, 2020; Birhane et al., 2022b; Scheuerman et al., 2021; Hutchinson et al., 2022; Bommasani, 2022; Denton et al., 2020, 2021; Mathur et al., 2022; Shilton, 2018; Broussard et al., 2019; Green, 2021; Blodgett et al., 2020; Viljoen, 2021; Abebe et al., 2020; Birhane and Guest, 2021; Blili-Hamelin and Hancox-Li, 2023; Costanza-Chock, 2020). AGI and human-level AI concern not only existing technologies, but also the technologies that many AI builders, researchers, and organizations dream of building in the future. When companies describe their official “long term aim [as] to solve intelligence, developing more general and capable problem-solving systems, known as artificial general intelligence (AGI)” (DeepMind, 2022), or when thought leaders claim that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” (Center for AI Safety, [n. d.]), they engage in “the process of negotiating between competing perspectives, values, and goals” (Green, 2021).
@@
We interrogate the political, social, and ethical questions undergirding AGI discourse by paying attention to how different accounts embody competing visions for the future of technology. We don’t see value-laden assumptions as a flaw. We see them as choices that admit legitimate disagreement through competing values. Our account of alternative paths forward (Section 4) strives to be reflective and explicit about the political and social assumptions of the visions we call for—such as democracy, epistemic justice, contextualism, and participation. Throughout, we also embrace the value of reflectiveness (Boyarskaya et al., 2020; Prunkl et al., 2021) about political, social, and ethical assumptions. We believe that making value assumptions explicit to ourselves and to others often makes for better individual and collective decisions—we might say, for more intelligent “experiments in living” (Mill, 2003; Anderson, 1991). Our aim in emphasizing reflectiveness is to point to more legitimate methods of determining the values that we want in AI. When large corporations are the only entities championing the narrative that their visions of AGI will transform everyone’s lives, what and whose values ultimately win out becomes a question of power divorced from legitimacy (Burrell and Metcalf, 2024).
@@
2.1. Intelligence is value-laden because it is a thick concept
@@
Intelligence is what philosophers would call a thick evaluative concept: it includes both descriptive and normative elements (Anderson, 2002; Kirchin, 2013, 2017). It contains descriptive elements about what empirical phenomena fall under the concept of intelligent behavior. It also contains a normative element. Evaluating intelligence inevitably involves assessing the desirability of specific behaviors. Previous research (Anderson, 2002; Alexandrova and Fabian, 2022; Blili-Hamelin and Hancox-Li, 2023; Cave, 2020) has argued that when we try to define or measure thick concepts, we inevitably embed ethical values in the design decisions we make about the boundaries of the concepts and the measurement methodology. Other examples of thick concepts that have value-laden definitions are health and well-being (Alexandrova and Fabian, 2022). The fact that intelligence is value-laden is at the root of debates over its definition and validity.
@@
The validity of the concept of general intelligence in humans has been extensively criticized. One of the principle targets of these critiques is Spearman’s g: a hidden, not directly observable factor expressing “shared variance across a set of intercorrelating cognitive tasks” (Warne and Burningham, 2019). Here we recap a few of the main critiques to provide some background for our later points about how some definitions of AGI make moves that are subject to similar critiques.
2.1.1. History of ableism and racism
@@
General intelligence metrics for humans have been criticized as ableist and racist (Anderson, 2002; Summerfield, 2023). Anti-racist critiques of human intelligence tests cite the influence of teacher expectations and stereotype threat, which affect test-takers of different demographics differently (Osborne, 2001). Disability advocates have argued that conventional standardized tests for intelligence are ableist because they do not consider how disabled individuals can perform within a broader social context (Pellicano and den Houting, 2022). In general, these critiques of g draw on the fact that human experiences and behaviors are diverse and contextual in ways that are hard to capture in standardized tests (Legault et al., 2021).
2.1.2. Weaknesses of factor analysis
@@
The use of factor analysis to discover g as a causal factor has been extensively critiqued (Glymour, 1998; Johnson, 2016). These critics argue that factor analysis, on its own, does not rule out alternative causal structures that could explain the same empirical results. Factor analysis can show that a common cause is a possible explanation of observed statistical patterns. But those same patterns could be explained by many other causal hypotheses.
@@
We mention this critique of factor analysis not just to situate the epistemic status of human intelligence metrics, but also as background for our later discussion (Section 3.2.9) of attempts in AI to find the equivalent of g in AI systems. These attempts are subject to the same critique that factor analysis does not rule out alternative causal hypotheses.
@@
2.1.3. Circularity
Another critique of psychometrics work purporting to quantify human intelligence is that it cannot be validated without reference to intelligence tests, making such measurements circular (Summerfield, 2023; Boring, 1923; Richardson, 2017; Popper, 2013). As Boring (1923) put it over 100 years ago, “[I]ntelligence is what the tests test”. Boring (1923)’s worry is not simply that intelligence tests need to be appealed to in validating intelligence tests. Rather, it is that the theoretical construct they purport to measure is itself motivated by setting aside the broad range of connotations of intelligence in favor of focusing on whatever can be measured through intelligence tests (Boring, 1923). This critique points to how psychometrics research is insufficiently motivated and under-theorized. Warne and Burningham (2019) is an example of recent work that doubles down on disconnecting research on g from any attempt at anchoring psychometrics research in a substantive theoretical conception of intelligence. As with the critique of factor analysis (2.1.2), deflationary approaches that avoid ascribing an explanatory role to g may be a potential response to this worry.  [^1] For an overview of recent approaches to motivating psychomterics research, see Deary (2012).
2.2. Relevance to AGI
@@
3. The motley choices of AGI discourse
@@
In this section, we look at some of the typical moves made when defining AGI. Our central contention is that these definitions always depend on assumptions about what is valuable to some group of people or what goals are worthwhile—whether implicitly or explicitly. This is expected once we understand that the topic of AGI and human-level AI sits at the intersection of two fundamentally value-laden questions. The topic of intelligence is fundamentally value-laden due to its nature as a thick concept (see Section 2). Moreover, the topic of what technologies and tools are worth building is itself fundamentally value-laden (Costanza-Chock, 2020; Birhane et al., 2022b; Dotan and Milli, 2020). No conception of AGI can escape the political, social, and ethical priorities that do and should shape any answer to the question of what is worth building, by whom, and for what purposes. Throughout this section, we highlight value-laden aspects of AGI discourse that pertain both to its character as a technology and to the nature of the task of defining intelligence.
@@
Some accounts of AGI attempt to adopt a kind of neutral stance. They may start by declaring that they intend to have a deflationary, value-neutral take on AGI (similar to the deflationary view on g we sketched in Section 2.1.2), but even these ultimately end up going beyond a purely deflationary view (see Section 3.3). We start by sketching the deflationary view (Section 3.1), then proceed to look at some of the major AGI definitions out there and identify dimensions on which different definitions make different value-laden assumptions (Section 3.2).
3.1. Purportedly deflationary accounts of AGI
@@
Some definitions of AGI start out with seemingly neutral goals, such as creating a shared language or common standards for measuring AGI, without (at least at first) specifying any normative goals beyond that. This can be viewed as similar to the deflationary view of g described in Section 2.1.2: where g is simply a correlational factor that you can measure from a battery of tests and does not necessarily represent any underlying causal or explanatory factor. One can adopt standards for the sake of being able to consistently compare systems against one another within those standards without assuming that those standards are the only “real” or “objective” standards.
@@
An example of the deflationary view is represented (at least initially) in “Levels of AGI” (Morris et al., 2023). At the start of the paper, the authors claim that they are seeking a “common language to compare models, assess risks, and measure progress along the path to AGI.” They think that having a “clear, operational definition of AGI” is the way to do this.
@@
At first, their goal seems to be purely deflationary—they want some shared, operationalizable terminology so that we know when we are referring to similar phenomena. Morris et al. (2023) are inspired by the Levels of Automation framework (SAE International, 2021) used to grade self-driving car technology—a framework that was also allegedly adopted to enhance communication. [^2] However, see Section 3.3 for a discussion of how this framework also drifted from its initial deflationary goals. Their deflationary approach is further underlined by a willingness to change the definition of AGI in the future if we gain the ability to operationalize things that are currently not operationalizable. For example, they currently want to define AGI in terms of “capabilities rather than processes” because we currently have little insight into the underlying mechanisms of many AI systems. They allow that research into mechanistic interpretability might eventually lead to operationalizable definitions involving processes, which then “may be relevant to future definitions of AGI” (Morris et al., 2023).
@@
Another AGI account that is on the deflationary spectrum is that of Adams et al. (2012), who have “a pragmatic goal for measuring progress toward its attainment”. The authors acknowledge that “The heterogeneity of general intelligence in humans makes it practically impossible to develop a comprehensive, fine-grained measurement system for AGI”. However, in order to measure progress, they would still like “a common framework for collaboration and comparison of results”. This conception is similar to that of (Morris et al., 2023) in treating the AGI framework as a common language for communication purposes. To the extent that Adams et al. (2012) adopt requirements for “general cognitive architectures”, they emphasize that these requirements are not final and are “simply …a convenient point of departure for discussion and collaboration”.
@@
3.2. Value-laden choices in AGI definitions: a taxonomy

### v4

@@ Abstract
1 Introduction
There is no agreed-upon definition of artificial general intelligence (AGI) (Morris et al. 2023; Mitchell 2024). Yet from influential AI companies (DeepMind 2022; OpenAI 2018; Meaker 2023; Ingram 2024) and AI researchers (Agüera y Arcas and Norvig 2023; OpenAI 2018; Morris et al. 2023; Chollet 2019; Bubeck et al. 2023) to the increasingly public worry (Future of Life Institute 2023; Center for AI Safety 2024) about existential risks (Bostrom 2014; McLean et al. 2023; Naudé and Barten 2023; Noy and Uher 2022), AGI and human-level AI have become one of the dominant ways to imagine the future potential of AI. At the same time, AGI has had its own skeptics (LeCun 2022; Toews 2022; Heaven 2020; Gebru and Torres 2024) who risk taking the target of their critiques to be more homogenous and congruent than it actually is.
The lack of homogeneity in current conceptions of AGI is not a bug. It is a feature of the underlying topic: what might it mean for machines to have intelligence that rivals human intelligence? Substantive disagreements come with the value-laden character of both intelligence and of AI as a technology. By value-laden, we mean that political, social, and ethical values can and should shape conceptions of intelligence, technology, and their intersection.(Blili-Hamelin and Hancox-Li 2023) We lack consensus on the definition of intelligence because of conflicting values. The same applies to the current lack of consensus on AGI.
Ruha Benjamin warns that “a narrow definition of what even counts as technology or intelligence” threatens the ability of communities to imagine worlds worth building.(Benjamin 2024) What forms of intelligence and technology are worth imagining are political and social questions. However, current approaches to AGI risk mistaking these questions for technical questions. We lay the groundwork for examining these potential mistakes by analyzing and classifying contingent assumptions underlying different definitions of AGI. Contending with the fractured worldviews among conceptions of AGI plants the seeds for resisting harmfully narrow conceptions of machine intelligence.
@@ 1 Introduction
Our investigation is not limited to accounts using the exact phrase “artificial general intelligence”. Like Morris et al. 2023, our topic is the long history of treating “[a]chieving human-level “intelligence” as the “north-star goal” of the AI field (Morris et al. 2023), dating at least as far back as the 1955 Dartmouth AI Conference (McCarthy et al. 1955). Discussions of “human-level AI”, “general AI”, or “AI” (such as in the phrase “strong AI” (Searle 1980)) fall within the scope of our account.
@@
2 Between human intelligence and technology: AGI’s dual value-laden pedigrees
@@
Building on STS scholarship on the values embedded in technology (Winner 1980), research communities like FAccT, AIES, and CHI have taken a deep interest in the political, social, and ethical values embedded both in AI tools and in the practices that surround AI (Fishman and Hancox-Li 2022; Dotan and Milli 2020; Birhane et al. 2022b; Scheuerman, Hanna, and Denton 2021; Hutchinson et al. 2022; Bommasani 2022; Denton et al. 2020, 2021; Mathur, Lustig, and Kaziunas 2022; Shilton 2018; Broussard et al. 2019; Green 2021; Blodgett et al. 2020; Viljoen 2021; Abebe et al. 2020; Birhane and Guest 2021; Blili-Hamelin and Hancox-Li 2023; Costanza-Chock 2020). AGI and human-level AI concern not only existing technologies, but also the technologies that many AI builders, researchers, and organizations dream of building in the future. When companies describe their official “long term aim [as] to solve intelligence, developing more general and capable problem-solving systems, known as artificial general intelligence (AGI)” (DeepMind 2022), or when thought leaders claim that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” (Center for AI Safety 2024), they engage in “the process of negotiating between competing perspectives, values, and goals” (Green 2021).
@@
We interrogate the political, social, and ethical questions undergirding AGI discourse by paying attention to how different accounts embody competing visions for the future of technology. We don’t see value-laden assumptions as a flaw. We see them as choices that admit legitimate disagreement through competing values. Our account of alternative paths forward (Section 4) strives to be reflective and explicit about the political and social assumptions of the visions we call for—such as democracy, epistemic justice, contextualism, and participation. Throughout, we also embrace the value of reflectiveness (Boyarskaya, Olteanu, and Crawford 2020; Prunkl et al. 2021) about political, social, and ethical assumptions. We believe that making value assumptions explicit to ourselves and to others often makes for better individual and collective decisions—we might say, for more intelligent “experiments in living” (Mill 2003; Anderson 1991). Our aim in emphasizing reflectiveness is to point to more legitimate methods of determining the values that we want in AI. When large corporations are the only entities championing the narrative that their visions of AGI will transform everyone’s lives, what and whose values ultimately win out becomes a question of power divorced from legitimacy (Burrell and Metcalf 2024).
@@
2.1 Intelligence is value-laden because it is a thick concept
@@
Intelligence is what philosophers would call a thick evaluative concept: it includes both descriptive and normative elements (Anderson 2002; Kirchin 2013, 2017). It contains descriptive elements about what empirical phenomena fall under the concept of intelligent behavior. It also contains a normative element. Evaluating intelligence inevitably involves assessing the desirability of specific behaviors. Previous research (Anderson 2002; Alexandrova and Fabian 2022; Blili-Hamelin and Hancox-Li 2023; Cave 2020) has argued that when we try to define or measure thick concepts, we inevitably embed ethical values in the design decisions we make about the boundaries of the concepts and the measurement methodology. Other examples of thick concepts that have value-laden definitions are health and well-being (Alexandrova and Fabian 2022). The fact that intelligence is value-laden is at the root of debates over its definition and validity.
@@
The validity of the concept of general intelligence in humans has been extensively criticized. One of the principle targets of these critiques is Spearman’s g: a hidden, not directly observable factor expressing “shared variance across a set of intercorrelating cognitive tasks” (Warne and Burningham 2019). Here we recap a few of the main critiques to provide some background for our later points about how some definitions of AGI make moves that are subject to similar critiques.
History of ableism and racism
@@
General intelligence metrics for humans have been criticized as ableist and racist (Anderson 2002; Summerfield 2023). Anti-racist critiques of human intelligence tests cite the influence of teacher expectations and stereotype threat, which affect test-takers of different demographics differently (Osborne 2001). Disability advocates have argued that conventional standardized tests for intelligence are ableist because they do not consider how disabled individuals can perform within a broader social context (Pellicano and den Houting 2022). In general, these critiques of g draw on the fact that human experiences and behaviors are diverse and contextual in ways that are hard to capture in standardized tests (Legault, Bourdon, and Poirier 2021).
Weaknesses of factor analysis
@@
The use of factor analysis to discover g as a causal factor has been extensively critiqued (Glymour 1998; Johnson 2016). These critics argue that factor analysis, on its own, does not rule out alternative causal structures that could explain the same empirical results. Factor analysis can show that a common cause is a possible explanation of observed statistical patterns. But those same patterns could be explained by many other causal hypotheses.
@@
We mention this critique of factor analysis not just to situate the epistemic status of human intelligence metrics, but also as background for our later discussion (Section 3.2) of attempts in AI to find the equivalent of g in AI systems. These attempts are subject to the same critique that factor analysis does not rule out alternative causal hypotheses.
@@
Circularity
Another critique of psychometrics work purporting to quantify human intelligence is that it cannot be validated without reference to intelligence tests, making such measurements circular (Summerfield 2023; Boring 1923; Richardson 2017; Popper 2013). As Boring (1923) put it over 100 years ago, “[I]ntelligence is what the tests test”. Boring (1923)’s worry is not simply that intelligence tests need to be appealed to in validating intelligence tests. Rather, it is that the theoretical construct they purport to measure is itself motivated by setting aside the broad range of connotations of intelligence in favor of focusing on whatever can be measured through intelligence tests (Boring 1923). This critique points to how psychometrics research is insufficiently motivated and under-theorized. Warne and Burningham (2019) is an example of recent work that doubles down on disconnecting research on g from any attempt at anchoring psychometrics research in a substantive theoretical conception of intelligence. As with the critique of factor analysis (2.1), deflationary approaches that avoid ascribing an explanatory role to g may be a potential response to this worry.  [^1] For an overview of recent approaches to motivating psychomterics research, see Deary (2012).
2.2 Relevance to AGI
@@
3 The motley choices of AGI discourse
@@
In this section, we look at some of the typical moves made when defining AGI. Our central contention is that these definitions always depend on assumptions about what is valuable to some group of people or what goals are worthwhile—whether implicitly or explicitly. This is expected once we understand that the topic of AGI and human-level AI sits at the intersection of two fundamentally value-laden questions. The topic of intelligence is fundamentally value-laden due to its nature as a thick concept (see Section 2). Moreover, the topic of what technologies and tools are worth building is itself fundamentally value-laden (Costanza-Chock 2020; Birhane et al. 2022b; Dotan and Milli 2020). No conception of AGI can escape the political, social, and ethical priorities that do and should shape any answer to the question of what is worth building, by whom, and for what purposes. Throughout this section, we highlight value-laden aspects of AGI discourse that pertain both to its character as a technology and to the nature of the task of defining intelligence.
@@
Some accounts of AGI attempt to adopt a kind of neutral stance. They may start by declaring that they intend to have a deflationary, value-neutral take on AGI (similar to the deflationary view on g we sketched in Section 2.1), but even these ultimately end up going beyond a purely deflationary view (see Section 3.3). We start by sketching the deflationary view (Section 3.1), then proceed to look at some of the major AGI definitions out there and identify dimensions on which different definitions make different value-laden assumptions (Section 3.2).
3.1 Purportedly deflationary accounts of AGI
@@
Some definitions of AGI start out with seemingly neutral goals, such as creating a shared language or common standards for measuring AGI, without (at least at first) specifying any normative goals beyond that. This can be viewed as similar to the deflationary view of g described in Section 2.1: where g is simply a correlational factor that you can measure from a battery of tests and does not necessarily represent any underlying causal or explanatory factor. One can adopt standards for the sake of being able to consistently compare systems against one another within those standards without assuming that those standards are the only “real” or “objective” standards.
@@
An example of the deflationary view is represented (at least initially) in “Levels of AGI” (Morris et al. 2023). At the start of the paper, the authors claim that they are seeking a “common language to compare models, assess risks, and measure progress along the path to AGI.” They think that having a “clear, operational definition of AGI” is the way to do this.
@@
At first, their goal seems to be purely deflationary—they want some shared, operationalizable terminology so that we know when we are referring to similar phenomena. Morris et al. (2023) are inspired by the Levels of Automation framework (SAE International 2021) used to grade self-driving car technology—a framework that was also allegedly adopted to enhance communication. [^2] However, see Section 3.3 for a discussion of how this framework also drifted from its initial deflationary goals. Their deflationary approach is further underlined by a willingness to change the definition of AGI in the future if we gain the ability to operationalize things that are currently not operationalizable. For example, they currently want to define AGI in terms of “capabilities rather than processes” because we currently have little insight into the underlying mechanisms of many AI systems. They allow that research into mechanistic interpretability might eventually lead to operationalizable definitions involving processes, which then “may be relevant to future definitions of AGI” (Morris et al. 2023).
@@
Another AGI account that is on the deflationary spectrum is that of Adams et al. (2012), who have “a pragmatic goal for measuring progress toward its attainment”. The authors acknowledge that “The heterogeneity of general intelligence in humans makes it practically impossible to develop a comprehensive, fine-grained measurement system for AGI”. However, in order to measure progress, they would still like “a common framework for collaboration and comparison of results”. This conception is similar to that of (Morris et al. 2023) in treating the AGI framework as a common language for communication purposes. To the extent that Adams et al. (2012) adopt requirements for “general cognitive architectures”, they emphasize that these requirements are not final and are “simply …a convenient point of departure for discussion and collaboration”.
@@
3.2 Value-laden choices in AGI definitions: a taxonomy